# DAO: A New and Public Non-Relativistic Reflection Model

### v1

@@ 3.6 Effects of varying the hydrogen density
4 discussion
In this paper, we presented a new reflection model, which we named DAO. This model is open-source and highly flexible for studying the reflection phenomenon in accretion-disk-corona systems. The model assumes a slab geometry, with the parameters of the slab and the incident radiation field listed in Table 1. It uses the XSTAR code to calculate the photoionization of each layer and the Feautrier method to solve the radiative transfer through the layers.
We compared our new X-ray reflection model DAO with xillver and reflionx. First, DAO includes a more complete and up-to-date atomic database than that of reflionx. Second, DAO implements a more accurate treatment of Compton scattering than the current xillver and reflionx tables (see Figures 3 and 4). Third, DAO is open-source and allows users to freely define their own incident spectrum. This flexibility is crucial for several applications—for example, modeling returning radiation (2024ApJ...976..229M; 2024ApJ...965...66M), or describing reflection in Z sources where the illuminating spectrum is neither a pure power law nor a blackbody, but instead a saturated Comptonization continuum (Ludlam2024Ap&SS.369...16L).
In the future, we plan to extend our model to higher hydrogen densities to better represent environments with dense gas in XRBs and AGNs (e.g., 2019MNRAS.489.3436J; 2023ApJ...951..145L), and to incorporate disk-blackbody radiation from the bottom boundary, which is particularly important for X-ray binaries (e.g., 2007MNRAS.381.1697R). We will also implement a non-uniform vertical density structure and account for hydrostatic equilibrium (2007MNRAS.377L..59D). Another major development will be the inclusion of polarization in the radiative transfer calculations, including Compton scattering for polarized X-ray reflection spectra. We further intend to examine the impact of different atomic databases and photoionization codes, for example by integrating CLOUDY (Gunasekera2025arXiv250801102G) into our framework.
@@ Appendix A Algorithm
where $\mathbf{T}$ represents the coefficient matrix, $\mathbf{u}$ is the vector of mean intensities, and $\mathbf{R}$ is the corresponding source vector. Equation A9, together with the boundary condition in Eq. A7, is solvable via a recursive forward-elimination and back-substitution procedure (1964CR....258.3189F; 1967AnAp...30..125F; 1968AnAp...31..257F). This solution scheme is collectively known as the Feautrier method.
We validate our radiative transfer code by comparing its results with the established reflection model pexrav (1995MNRAS.273..837M), as illustrated in Figure 17. The spectra presented in Figure 17 demonstrate excellent agreement in the continuum below 10keV10\ $\mathrm${keV}. While the pexrav spectrum exhibits a slightly higher peak in the Compton Hump range, the two models agree well at energies E<20keVE<20\ $\mathrm${keV} and E>50keVE>50\ $\mathrm${keV}. We hypothesize that this difference will decrease upon the inclusion of the angle-dependent source function, as described in Section 2. It must be noted that the pexrav model does not incorporate line emission; however, this exclusion is negligible when focusing on the high-energy scattered continuum.
@@ B.1 Gaussian approximation redistribution function
The Gaussian redistribution function was first applied to the scattering problem in Compton-thick atmospheres by 1978ApJ...219..292R; 1993MNRAS.261...74R to describe the down-scattering of photons with $E<200$ keV, and was later extended by 2000ApJ...537..833N to treat Compton scattering over the full energy range.
@@ B.2 Exact Redistribution Function
We adopt the exact redistribution function following the formalism of 1993AstL...19..262N and 2000ApJ...537..833N:
@@
$R(x,x_{1},\mu)=\frac{3}{32\mu\Theta K_{2}(1/\Theta)}\int_{\left(x-x_{1}+Q\sqrt{1+2/q}\right)/2}^{\infty}R_{E}(x,x_{1},\mu,\gamma)\exp(-\gamma/\Theta)\,d\gamma,$ (B7)
@@
$R(x,x_{1})=\int_{\mu_{\mathrm{min}}}^{\mu_{\mathrm{max}}}R(x,x_{1},\mu)\,d\mu.$ (B8)
Figure 18 illustrates the significant deviation between the exact redistribution function and the Gaussian approximation when the gas temperature is high and the photon energies are large. Similar results were reported by 2020ApJ...897...67G; the comparison is reproduced here to facilitate the reader’s understanding of the differences. The redistribution function in Eq. B8 satisfies the normalization condition with respect to the Compton scattering cross section:
@@
A detailed comparison between $\sigma_{\mathrm{cs}}$ and $\sigma_{\mathrm{KN}}$ is provided in 2020ApJ...897...67G.
@@ Appendix C Relativistic spectrum
The reldao and reldaoA models are developed within the framework of relxill v2.5 (2014ApJ...782...76G; 2025ApJ...989..168H) and relxillA (2025ApJ...989..168H). Currently, as the full table models have not yet been generated, reldao serves as a prototype rather than a finalized relativistic reflection model. The code will be publicly released upon the completion of the DAO table models. Additionally, the relativistic spectrum is computed using the convolution model relconv (2010MNRAS.409.1534D) within XSPEC (1996ASPC..101...17A), yielding the combined model relconv*dao.
@@
All the physical parameters above have been explained in detail in section 2 of 2025ApJ...989..168H.

### v2

@@ III.6 Effects of varying the hydrogen density
IV discussion
In this paper, we presented a new reflection model, which we named DAO. This model is open-source and highly flexible for studying the reflection phenomenon in accretion-disk-corona systems. The model assumes a slab geometry, with the parameters of the slab and the incident radiation field listed in Table 1. Following the mathematical framework of the xillver models, DAO uses the XSTAR code to calculate the photoionization of each layer and the Feautrier method to solve the radiative transfer through the layers.
We compared our new X-ray reflection model DAO with xillver and reflionx. Both DAO and xillver include more complete atomic databases, and a better treatment of angular dependence, than reflionx. DAO includes several features that were recently included in the proprietary xillver code, but have not yet been released as publicly available xillver table models. First, DAO includes the most complete and up-to-date atomic database available. This was explored for the xillver model by Ding et al. (2024) but has not yet been included in a publicly available table model. DAO implements an accurate treatment of Compton scattering, which was explored for the xillver model by García et al. (2020) but again has not yet been publicly released in a table model. We also include some effects that, to our knowledge, are not considered in xillver, such as the inclusion of nickel and employing a slightly more advanced treatment of the incidence angle. However, we believe that the main advantage of DAO is that it is open-source and allows users to freely define their own incident spectrum. This flexibility is crucial for several applications—for example, modelling returning radiation (Mirzaev et al. 2024a, b), or describing reflection in Z sources where the illuminating spectrum is neither a pure power law nor a blackbody, but instead a saturated Comptonization continuum (Ludlam 2024).
In the future, we plan to extend our model to higher hydrogen densities to better represent environments with dense gas in XRBs and AGNs (e.g., Jiang et al. 2019; Liu et al. 2023b), and to incorporate disk-blackbody radiation from the bottom boundary, which is particularly important for X-ray binaries (e.g., Ross and Fabian 2007). We will also implement a non-uniform vertical density structure and account for hydrostatic equilibrium (Done and Nayakshin 2007). Another major development will be the inclusion of polarization in the radiative transfer calculations, including Compton scattering for polarized X-ray reflection spectra. We further intend to examine the impact of different atomic databases and photoionization codes, for example by integrating CLOUDY (Gunasekera et al. 2025) into our framework.
@@ Appendix A Algorithm
where $\mathbf{T}$ represents the coefficient matrix, $\mathbf{u}$ is the vector of mean intensities, and $\mathbf{R}$ is the corresponding source vector. Equation A9, together with the boundary condition in Eq. A7, is solvable via a recursive forward-elimination and back-substitution procedure (Feautrier 1964, 1967, 1968). This solution scheme is collectively known as the Feautrier method.
We validate our radiative transfer code by comparing its results with the established reflection model pexrav (Magdziarz and Zdziarski 1995), as illustrated in Figure 17. The spectra presented in Figure 17 demonstrate excellent agreement in the continuum below 10keV10\ $\mathrm${keV}. While the pexrav spectrum exhibits a slightly higher peak in the Compton Hump range, the two models agree well at energies E<20keVE<20\ $\mathrm${keV} and E>50keVE>50\ $\mathrm${keV}. We hypothesize that this difference will decrease upon the inclusion of the angle-dependent source function, as described in Section II. It must be noted that the pexrav model does not incorporate line emission; however, this exclusion is negligible when focusing on the high-energy scattered continuum.
@@ B.1 Gaussian approximation redistribution function
The Gaussian redistribution function was first applied to the scattering problem in Compton-thick atmospheres by Ross et al. (1978); Ross and Fabian (1993) to describe the down-scattering of photons with $E<200$ keV, and was later extended by Nayakshin et al. (2000) to treat Compton scattering over the full energy range.
@@ B.2 Exact Redistribution Function
Following the study of García et al. (2020), we use the method described in Nagirner and Poutanen (1993); Madej et al. (2017) to calculate the exact redistribution function
@@
$R(x,x_{1},\mu)=\frac{3}{32\mu\Theta K_{2}(1/\Theta)}\int_{\left(x-x_{1}+Q\sqrt{1+2/q}\right)/2}^{\infty}R_{E}(x,x_{1},\mu,\gamma)\exp(-\gamma/\Theta)\,d\gamma$ (B7)
@@
$R(x,x_{1})=\int_{\mu_{\mathrm{min}}}^{\mu_{\mathrm{max}}}R(x,x_{1},\mu)\,d\mu$ (B8)
Figure 18 illustrates the significant deviation between the exact redistribution function and the Gaussian approximation when the gas temperature is high and the photon energies are large. Similar results were reported by García et al. (2020); the comparison is reproduced here to facilitate the reader’s understanding of the differences. The redistribution function in Eq. B8 satisfies the normalization condition with respect to the Compton scattering cross section:
@@
A detailed comparison between $\sigma_{\mathrm{cs}}$ and $\sigma_{\mathrm{KN}}$ is provided in García et al. (2020).
@@ Appendix C Relativistic spectrum
The reldao and reldaoA models are developed within the framework of relxill v2.5 (García et al. 2014; Huang et al. 2025b) and relxillA (Huang et al. 2025b). Currently, as the full table models have not yet been generated, reldao serves as a prototype rather than a finalized relativistic reflection model. The code will be publicly released upon the completion of the DAO table models. Additionally, the relativistic spectrum is computed using the convolution model relconv (Dauser et al. 2010) within XSPEC (Arnaud 1996), yielding the combined model relconv*dao.
@@
All the physical parameters above have been explained in detail in section 2 of Huang et al. (2025b).